OpenStudy (anonymous):

Hannah earned $1536 in 8 years on an investment at a 4% annual simple interest rate. how much was hannah's investment?

OpenStudy (agl202):

4% a year with simple interest for 8 years = 8 * 4% = 32%. 32% of Investment = 1536. Investment = 1536/32% = 1536 * 100 / 32 = 4,800